rescue crews working for stockton rescue average 3 calls per shift. after his first four shifts, brian had recieved 5,2,1, and 3 calls. how many calls will brian need on his next shift
Answer by rfer(12730)

5+2+1+3+x/5=3
11+x/5=3
11+x=15
x=15-11
x=4